# A Reanalysis of the FDA’s Benefit-Risk Assessment of Moderna’s mRNA-1273 COVID Vaccine: For 18–25-Year-Old Males, Risks Exceeded Benefits Relative to Hospitalizations

### v1

@@ 4. Reanalysis: Results
In Table 6 below, we compare the hospitalizations-prevented benefit of mRNA-1273 vaccination to corresponding VAM/P-hospitalizations risk through our Scenarios A–E, with Scenarios A–C being reanalyses of the corresponding FDA Scenarios 1–3.
{adjustwidth}
-1.2cm
Table 6. Scenarios projecting the hospitalizations-prevented benefit of mRNA-1273 vaccination of a hypothetical test population of 1 million males ages 18–25 compared to the estimated corresponding VAM/P hospitalization risk $H_{\rm VAM/P}$. For the FDA’s Scenarios 1–3, the values of $U$ and $V$ are determined by the FDA’s model described in Section 2.2, while, for our Scenarios A-E, the values of $U$ and $V$ determined .by our model described in Section 3. A green background for a scenario indicates benefit exceeds risk.
Scenario Description Assumptions* $U$ $V$ Hospitalizations Prevented Per Million Vaccinations $U-V$ $H_{\rm VAM/P}$ Benefit Risk Ratio
FDA
Scenario 1
Highest 2021 COVID in- cidence assumed. FDA’s “most likely scenario”
@@
6619 1853 4766 110 (2nd-dose only) 43.33
A Reanalysis of FDA’s Scenario 1 $I_{r}=1,H_{r}=0.0014,\\ F_{pi}=0.62,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 629 168 461 268 1.72
FDA
Scenario 2
Average 2021 COVID incidence assumed
$U=2900$
$V=2900(1-0.72)\approx 812$
2900 812 2088 110 (2nd-dose only) 18.98
B Reanalysis of FDA’s Scenario 2 $I_{r}=0.356,H_{r}=0.0014,\\ F_{pi}=0.62,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 224 60 164 268 0.61
FDA
Scenario 3
  Lowest incidence of pandemic assumed  (6/5/21) $U=882$ $V=882(1-0.72)\approx 247$ 882 247 635 110 (2nd-dose only) 5.77
C Reanalysis of FDA’s Scenario 3 $I_{r}=0.052,H_{r}=0.0014,\\ F_{pi}=0.62,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 33 9 24 268 0.09
D COVID incidence   twice that for the 2nd COVID wave in the U.S. Our “most likely scenario” $I_{r}=0.456,H_{r}=0.0014,\\ F_{pi}=0.69,E_{pi}=0.47,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.79,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 247 68 179 268 0.67
E Equivalent protection for prior infected and COVID-naïve vaccinated; incidence as in Scenario D. $I_{r}=0.456,H_{r}=0.0014,\\ F_{pi}=0.69,E_{pi}=0.30,\\ E_{v}=0.30,E_{h}=0.57,\\ HRR_{pi}=0.67,HRR_{v}=0.67,\\ HRR_{h}=0.88$ 300 68 232 268 0.87
* For discussions of how “Assumptions” were derived, see Section 3.3 as well as Supplement 1.
To illustrate the application of our model, we carry out our Scenario-A computation. Using Scenario A assumptions for variable values and our definition (3.2.1) of the function $U$, we obtain
$U=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot 0.0014+0.62\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.47\right)\cdot 0.0014\cdot\left(1-0.79\right)\approx 629,\$ and, from our definition (3.2.2) of the function $V$,
$V=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.30\right)\cdot 0.0014\cdot\left(1-0.67\right)+0.62\cdot 1\cdot\left(1-0.57\right)\cdot 0.0014\cdot\left(1-0.88\right)\approx 168.$
@@
Observe that our Scenario-D projection of 179 COVID-hospital admissions prevented by vaccination corresponds to $43\approx 0.24\cdot 179$ ICU admissions prevented [9, Table 3], which should be compared to the 47 ICU admissions for VAM/P projected by the BLA-Memo Scenario and 0 such admissions projected in the scenarios of [9].
In Supplement 2, we carry out sensitivity analyses exploring how the benefit-risk picture for mRNA-1273 vaccination changes with modified assumptions for key variables. We also show that in Scenarios B–E, when hospitalizations prevented by vaccination include those prevented by dose 1 during the period preceding the point of full vaccination, vaccine risks still outweigh benefits in terms of the hospitalizations endpoint.
@@ 4.1. Subpopulation Analyses Based on Prior-Infection Status
The FDA acknowledges, “The health condition of individuals may remarkably impact the B-R profile and its evaluation” [9, Section 4]; however, the authors of [9] do not explore this remarkable impact, explaining “[W]e did not investigate the benefits and risks for specific subpopulations, such as those with comorbidities, due to limited information for these populations.” Comorbidities such as obesity may triple COVID-hospitalization risk [55]. In this subsection, we consider subpopulations for which there was sufficient information before January 2022 to conduct conclusive benefit-risk assessments. We consider, in particular, subpopulations having had a COVID infection prior to the evaluation period 1/1/22 – 5/31/22.
@@
Prior infection not only decreases the hospitalization-prevented benefit of vaccination but also increases the risk of a severe adverse event after vaccination: see, e.g., [56, 57], with [56] reporting prior infection is “associated with an increased risk of severe side effects leading to hospital care (1.56 [risk ratio] (95%CI: 1.14–2.12)).” Such “severe side effects” possibly linked to mRNA COVID vaccination and potentially requiring hospitalization include anaphylaxis [58], Multisystem Inflammatory Syndrome [59, 60], Guillain Barré Syndrome [61], posterior orthostatic tachycardia syndrome (POTS) [62], retinal vascular occlusion [63], and cerebral venous sinus thrombosis (CVST) [64]. Moreover, there is evidence that prior infection increases the risk of vaccine-associated myocarditis [43]—see Appendix 4 of our Supplement 2 for details. [^19] Survey results reported in [65] also suggest that prior infection may increase the risk of VAM/P; e.g., in Table 4, we find that 2.53% of those having a “laboratory-confirmed SARS-CoV-2 infection” before receiving dose 1 of mRNA-1273 reported “chest tightness/pain” during the 7-day period following the dose while only 0.54% of those not having a confirmed infection before dose 1 reported these symptoms; corresponding percentages for dose 2 (Table 5) are 1.78% and 0.98%.
Prior infection’s reducing the hospitalizations-prevented benefit of vaccination while increasing postvaccination adverse-event risk creates a compelling case for conducting a benefit-risk assessment for the subpopulation of unvaccinated 18–25-year-old males having prior-infection protection.
@@
Consider the subpopulation of 18–25-year-old unvaccinated males having a COVID-infection prior to the evaluation period 1/1/22 – 5/31/22. Using our model, we can compute the expected number of COVID hospitalizations for this subpopulation in various scenarios simply by setting $F_{pi}=1$ and leaving all other variables unchanged. For example, with $F_{pi}=1$ in Scenario A, our model predicts $U\approx 155.82$ hospitalizations per million and $V\approx 72.24$, yielding a hospitalizations-prevented benefit of 85.8 per million vaccinated, far below VAM/P hospitalization risk.
The analysis of the preceding paragraph may be reinterpreted in terms of evaluation-period hospitalization risk for a typical unvaccinated 18–25-year-old male, COVID-infected prior to the evaluation period. For ease of exposition, we will call this male Bob. [^20] If we were also considering COVID-hospitalization risk for a typical 18–25-year-old female, we would call her Alice, choosing “Alice” and “Bob” because these names are customarily used for “person A” and “person B” in the field of cryptography. If Bob chooses not to vaccinate, our model estimates the probability that Bob will be hospitalized for COVID during the evaluation period to be 155.82 in 1 million (assuming $I_{r}=1$). Bob’s total hospitalization risk associated with the choice to vaccinate is essentially the sum of his probability of VAM/P hospitalization and the probability of his having a breakthrough-case COVID hospitalization during the evaluation period (assuming postvaccination VAM/P hospitalization is independent of postvaccination COVID-hospitalization).212121 Technically speaking, because there is a small chance—on the order of 0.02 in 1 million—that Bob might experience both a VAM/P hospitalization and a breakthrough-COVID hospitalization, Bob’s total hospitalization risk from vaccination is the sum of his VAM/P-related risk and his COVID-related risk minus the product of these two risks (with the product being the small chance that Bob experiences both a VAM/P hospitalization and a COVID hospitalization by our assumption of independence of these events).
While there is some evidence that VAM/P risk increases with prior infection, we assume only that VAM/P risk does not decrease. Suppose Bob decides to take dose 1 and at the same time decides to take dose 2 if VAM/P does not occur after dose 1; then, if $r_{1}$, $r_{2}$ are, respectively, the VAM/P-incidence rates for doses 1 and 2 for males 18–25, then the probability Bob will experience VAM/P is well approximated by $r_{1}+(1-r_{1})r_{2}.$ Because $r_{1}$ is quite small, the preceding probability is essentially $r_{1}+r_{2}$, which explains why, even on the individual level, the FDA should consider the sum of risks, postdose 1 and postdose 2 in assessing the VAM/P risk of full vaccination (as the agency did in its population-level analyses of BNT162b2 [8, Section 2.3.1]) and of mRNA-1273 in its BLA-Memo [46, Section 4.7]). Therefore 268/1,000,000 approximates the probability that Bob experiences hospitalization owing to VAM/P during the course of mRNA-1273 vaccination. The probability that he experiences COVID hospitalization owing to a breakthrough infection is 72.24 in 1 million, according to our model. Thus, his total hospitalization risk from vaccination is approximately 340.24 in 1 million. This is under the Scenario-A assumption that $I_{r}=1$, which we interpret to mean that evaluation-period infection is certain for an infection-naïve male 18–25, and Bob’s probability of infection is $I_{r}(1-E_{pi})=0.53$. The following Table shows how Bob’s hospitalization risk varies as his risk of infection varies.
{adjustwidth}
-1.2cm
@@
$1$ 155.82 in 1 million 72.24 in 1 million $2.18\approx\frac{268+72.24}{155.82}$
@@
**$I_{r}(1-E_{pi})$ is the probability the “typical male” (18–25 with prior-infection protection) will become COVID infected over the evaluation period if unvaccinated; e.g., 0.53 if $I_{r}=1$
In Section 2.3 of Supplement 2, based on seroprevalence and modeling information, we estimate that $I_{r}\approx 0.5$ (and is likely $<0.5$), that is, roughly 1/2 of unvaccinated, infection-naïve 18–25-year-old males in the U.S. became infected over the evaluation period (and Bob’s infection risk is $0.2651/2(1-E_{pi})=$ 0.265 because of his prior infection). Thus, according to Table 7, our modeling suggests that for our typical unvaccinated 18–25-year-old male Bob with prior-infection, the choice to vaccinate creates a hospitalization risk approximately 4 times that associated with the choice not to vaccinate. Moreover, if we account for the increased likelihood of postvaccination adverse-event hospitalizations for those with prior infection, then the risk ratios of Table 7 grow even larger. Also note that for every value of $I_{r}$ in Table 7, hospitalization risk owing to VAM/P alone exceeds COVID-hospitalization risk.
Applying the information in Table 7 is difficult because a “typical male” has an “average comorbidities level” and the meaning of this is unclear. However, if our typical unvaccinated male Bob with prior-infection protection also has no comorbidities, then his hospitalization-risk ratios should be larger than those of Table 7. (There is no evidence that being free of comorbidities decreases VAM/P risk; e.g., according to a 7/16/21 report of the chief science advisor of Canada “Patients presenting with this complication [vaccine-associated myo/pericarditis] typically have been healthy individuals.” [66]. [^22] See also [67, p. 383].)
Now assume that Bob is an unvaccinated 18–25-year-old male with prior infection without comorbidities. To establish conclusively that Bob’s hospitalization risk is greater with vaccination, we consider two bounding analyses, each overestimating the benefits of vaccination and underestimating the adverse-event risks associated with vaccination.
{adjustwidth}
0.3in0.3in Scenario I: To underestimate hospitalization risk associated with vaccination, we assume that vaccination is 100% effective in preventing breakthrough-infection hospitalization, and we obtain a lower bound on Bob’s VAM/P-hospitalization risk, by ignoring dose 1 VAM/P hospitalizations and using our estimate of 242 VAM/P hospitalizations per million second doses of mRNA-1273 among males 18–25 based on Ontario data (as described in Table 3). Thus, ignoring his first-dose VAM/P hospitalization risk, his hospitalization risk for postvaccination adverse events other than myo/pericarditis, and his likely increased risk for adverse-event hospitalization owing to his prior infection, we obtain an underestimate of Bob’s postvaccination hospitalization risk of 242 in 1 million (based on VAM/P alone).
To overestimate Bob’s COVID-hospitalization risk assuming he remains unvaccinated, we assume that evaluation-period infection is certain for an infection-naïve unvaccinated male ($I_{r}=1$) and that Bob’s being free of comorbidities does not decrease his COVID hospitalization risk which is 155.82 in 1 million according to Table 7.
We conclude that Bob is more likely—at least $1.55\approx\frac{242}{155.82}$ times as likely—to be hospitalized for VAM/P after choosing to vaccinate than to be hospitalized for COVID while remaining unvaccinated.
{adjustwidth}
0.3in0.3in Scenario II: If we assume no COVID-hospitalization risk-reduction from Bob’s being free of comorbidities, assume his VAM/P-hospitalization risk is only 110 in 1 million (the FDA’s assumption in its most likely scenario of [9]), ignore hospitalization risks for other postvaccination adverse events (likely increased owing to Bob’s prior infection), and assume every infection-naïve unvaccinated male becomes infected over the evaluation period, then we still conclude that Bob is more likely—at least $1.17\approx\frac{110+72.24}{155.82}$ times as likely—to be hospitalized for VAM/P or a breakthrough infection after choosing to vaccinate than to be hospitalized for COVID while remaining unvaccinated. Note that if we use the VAM/P hospitalization risk estimate of the FDA’s BLA-Memo Scenario the ratio 1.17 rises to $1.29\approx\frac{128+72.24}{155.82}$.
@@
To further elucidate the important role that prior infection plays in a careful benefit-risk assessment of mRNA-1273 vaccination, we note that for an 18–25-year-old male without prior COVID infection, hospitalization-risk ratios become favorable for vaccination:
{adjustwidth}
-1.4cm0cm
@@
Table 8. Hospitalization-risk ratios for a typical male 18—25 without prior-infection protection: risk of hospitalization after choosing mRNA-1273 vaccination (VAM/P or severe COVID-breakthrough infection*) divided by risk of COVID hospitalization after choosing not to vaccinate.
@@
$1$ 1400 in 1 million 323.4 in 1 million $0.42\approx\frac{268+323.4}{1400}$
@@
*COVID hospitalization risk is considered over a 5-month evaluation period
@@ 4.2. Subpopulation Analyses Based on Prior-Infection Status and BMI
For the remainder of the paper, we no longer confine to footnotes discussions and citations of data unavailable at the time of the FDA’s benefit-risk analysis of mRNA-1273.
We illustrate how comorbidities may impact a benefit-risk assessment, focusing on obesity because evidence suggests that Body Mass Index (BMI) is the single most potent modifier of COVID-19 risk for those under 60 years old [68, 69]. According to the CDC [55], “Having obesity [BMI $\geq 30$] may triple the risk of hospitalization due to a COVID-19 infection.” This is consistent with a finding that hospitalization risk for the obese is 3.2 times that for the non-obese for an adolescent study population (ages 12–17) [70, Table IV]. Gao et al. [71] provide additional support for this CDC estimate based on a cohort study of 6,910,695 residents of England (data from the period 1/24/20 – 4/30/20). Gao et al. found that for those 20–39 years old COVID hospitalization risk is minimized when BMI $\approx 23$ and that for $23\leq\text{BMI}\leq 44$ risk is multiplied by a factor of approximately 1.09 (1.08, 1.10) for each unit increase in BMI. Gao et al. “found no difference in effect estimates by sex” [71, Figure 2]. Applying Gao’s model (see Appendix 5 of Supplement 2), we obtain the following:
Table 9. Estimated COVID hospitalization hazard ratio (HR) for 20–39 year-olds relative to those with BMI $\approx 23$ based on a model by Gao et al. [71]
BMI Range COVID Hospitalization HR
23–24 ref.
25–29 1.49
30–34 2.29
35–39 3.52
40–44 5.41
Assuming that for 18–25-year-old males obesity triples COVID-hospitalization risk and that 20.5% of this population is obese [72], we find not being obese drops baseline COVID hospitalization risk $H_{r}=0.0014$ to approximately $0.71\cdot H_{r}$ (see Appendix 5 of Supplement 2). We assume this same factor determines the drop in hospitalization risk for unvaccinated males 18–25 with prior infection (IHR $=0.71\cdot H_{r}\cdot(1-E_{pi})(1-HRR_{pi})$).
We assume that not being obese for an otherwise typical vaccinated 18–25-year-old male decreases IHR by a greater amount: IHR $=0.36\cdot H_{r}(1-E_{v})(1-HRR_{v})$ based on case-hospitalization rates computed from the raw data of Supplementary Table A2 of [73].
We provide below a COVID hospitalization benefit-risk analysis derived from Gao et al.’s BMI-based model [71] described above. We assume that the BMI range 23–24 for 18–25-year-old males provides minimum COVID-hospitalization risk, which we estimate by multiplying the hospitalization risks appearing in column 2 of Tables 7 and 8 by the factor 0.71 and those in column 3 by the factor of 0.36. For a BMI-based hazard ratio HR from Table 9, we compute the corresponding effect on the hospitalization risks in column 2 of Tables 7 and 8 via the factor $\text{HR}\cdot 0.71$ and for those in column 3 via the factor $\text{HR}\cdot 0.36$.
Table 10. Hospitalization-risk ratios (risk if vaccinated (VAM/P & breakthrough)risk if $\left(\frac{\text{risk if vaccinated (VAM/P \& breakthrough)}}{\text{risk if unvaccinated}}\right)$ based on BMI for males 18—25. Ratios displayed as pairs: “with prior-infection protection, without prior-infection projection.”
@@
**Sample computation: $\frac{268+36.12\cdot 1.49\cdot 0.36}{77.91\cdot 1.49\cdot 0.71},\frac{268+161.7\cdot 1.49\cdot 0.36}{700\cdot 1.49\cdot 0.71}\approx 3.49,0.48$
Consistent with Table 8, the preceding table suggests that for 18–25-year-old males without prior-infection protection, vaccination reduces hospitalization risk for all BMI ranges except for the ideal range (BMI 23–24) in a low-infection-risk scenario. Similarly, consistent with Table 7, we see that for 18–25-year-old males with prior-infection protection, vaccination increases hospitalization risk except for the those with very high BMI (35–44) in a high-infection-risk scenario. In a medium infection-risk scenario ($I_{r}=0.5$), our modeling suggests that for the non-obese (BMI $<30$) having prior-infection protection the choice to vaccinate creates a hospitalization risk 3.49 to 5.08 times that associated with the choice not to vaccinate.
5. Validating the Model

### v3

@@ 4. Reanalysis: Results
In Table 6 below, we compare the hospitalizations-prevented benefit of mRNA-1273 vaccination to corresponding VAM/P-hospitalizations risk, with Scenarios A–C being reanalyses of the corresponding FDA Scenarios 1–3.
Table 6. Scenarios comparing the hospitalizations-prevented benefit of mRNA-1273 vaccination of a hypothetical test population of 1 million males 18–25 to the corresponding VAM/P hospitalization risk $H_{\rm VAM/P}$. For the FDA’s Scenarios 1–3, the values of $U$ and $V$ are determined by the FDA’s model described in Section 2.2, while, for Scenarios A-E, the values of $U$ and $V$ are determined .by our model described in Section 3. A green background indicates benefit exceeds risk.
Scenario Description Assumptions* $U$ $V$ Hospitalizations Prevented Per Million Vaccinations UVU-V $H_{\rm VAM/P}$ Benefit Risk Ratio
FDA Scenario 1 Highest 2021 COVID incidence assumed. FDA’s “most likely scenario”
@@
6619 1853 4766 110 (Dose 2 only) 43.33
@@
FDA Scenario 2 Average 2021 COVID incidence assumed $U$ = 2900 VV= 2900(1-0.72) $\approx$ 812 2900 812 2088 110 (Dose 2 only) 18.98
[…]
FDA Scenario 3 Lowest 2021 COVID incidence assumed $U$ = 882 VV= 882(1-0.72) $\approx$ 247 882 247 635 110 (Dose 2 only) 5.77
[…]
E Equivalent protection for prior infected and COVID-naïve vaccinated; incidence as in Scenario D. $I_{r}=0.456$, $H_{r}=0.0014$, $F_{pi}=0.69$, $E_{pi}=0.30$, $E_{v}=0.30$, $E_{h}=0.57$, $HRR_{pi}$$HRR_{v}=0.67$, $HRR_{h}=0.88$ 300 68 232 268 0.87
* For discussions of “Assumptions,” see Section 3.3 as well as Supplement 1.
We illustrate our model’s application by carrying out Scenario-A computations, Using Scenario A assumptions from Table 6 and our definitions (3.2.1) and (3.2.2) of $U$ and $V$, we obtain
$U=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot 0.0014+0.62\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.47\right)\cdot 0.0014\cdot\left(1-0.79\right)\approx 629,{\rm and}\\$ $V=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.30\right)\cdot 0.0014\cdot\left(1-0.67\right)+0.62\cdot 1\cdot\left(1-0.57\right)\cdot 0.0014\cdot\left(1-0.88\right)\approx 168.$
Thus, for Scenario A, the hospitalizations-prevented benefit of vaccination is projected to be UV= 629168=461U-V=\ 629-168=461 per million full mRNA-1273 vaccinations, roughly one-tenth of the FDA’s Scenario-1 estimate 4,766 [11]. Although Scenario A has a benefit-risk ratio above 1, its assumption that $I_{r}=1$ makes this scenario highly unlikely. [^9] Based on seroprevalence and modeling information, we estimate $I_{r}\leq 0.5$ [Supplement 2, Section 2.3]. For the more likely scenarios B, D, and E, our modeling suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from VAM/P alone compared to COVID hospitalizations prevented. Comparing the projections of FDA’s “most likely scenario” to ours, we find that the FDA’s benefit-risk ratio, 43.33, is over 60 times more favorable than ours, 0.67. Though Scenario D is our most likely scenario, Scenario E is equally important, being a “best-case scenario” for vaccination incorporating the minimum reasonable assumption of prior-infection protection vs. vaccine protection.
In Section 2 of Supplement 2, we carry out sensitivity analyses exploring how modifying model inputs impacts the benefit-risk picture. For example, if we fix our model-input values, except for $I_{r}$, at their Scenario-D levels, then benefit-risk equipoise is reached when $I_{r}\approx 0.685$. We also illustrate how slider-plots can provide insights concerning equipoise questions involving variation of more than one model input.
@@ 4.1. Subpopulation Analyses Based on Prior-Infection Status
There is a compelling case for conducting a benefit-risk assessment for the subpopulation of unvaccinated 18–25-year-old males having prior-infection protection because prior infection not only decreases the hospitalizations-prevented benefit of vaccination but also increases the risk of severe adverse events101010e.g., those described in [57, 58, 59, 60, 61, 62, 63] after vaccination [64, 65], with [64] reporting prior infection is “associated with an increased risk of severe side effects leading to hospital care (1.56 [risk ratio] (95%CI: 1.14–2.12)).” Moreover, prior infection may increase the risk of VAM/P [42]—see Appendix 4 of Supplement 2 for details. [^11] Survey results reported in [66] also suggest that prior infection may increase the risk of VAM/P; e.g., in Table 4, we see that 2.53% of those having a confirmed SARS-CoV-2 infection before dose 1 of mRNA-1273 reported “chest tightness/pain” during the 7-day period after the dose while only 0.54% of those not having a confirmed infection before dose 1 reported these symptoms; corresponding percentages for dose 2 (Table 5) are 1.78% and 0.98%.
@@
For the subpopulation of 18–25-year-old unvaccinated males having a COVID-infection prior to the evaluation period 1/1/22 – 5/31/22, we can use our model to compute the expected number of COVID hospitalizations in various scenarios simply by setting $F_{pi}=1$ while leaving all other inputs unchanged. For example, with $F_{pi}=1$ in Scenario A, our model predicts $U\approx 155.82$ hospitalizations per million and $V\approx 72.24$, yielding a hospitalizations-prevented benefit of 83.58 per million vaccinated, far below VAM/P hospitalization risk ($\approx$ 268 per million vaccinated).
As we discuss in Appendix 3 of Supplement 1, the preceding paragraph’s analysis may be reinterpreted in terms of evaluation-period hospitalization risk for a typical unvaccinated 18–25-year-old male, COVID-infected prior to the evaluation period. For such a male, the total hospitalization risk of the choice to vaccinate is essentially the sum of his probability of VAM/P hospitalization and his probability of having a breakthrough-case COVID hospitalization during the evaluation period. Our analysis is summarized in Tables 7 and 8 below in terms of hospitalization-risk ratios, with ratios <1<1 indicating the choice to vaccinate decreases risk and >1>1 indicating it increases risk.
@@
11 155.82 in 1 million 72.24 in 1 million $2.18\approx\frac{268+72.24}{155.82}$
@@
$I_{r}(1-E_{pi})$ is the probability the “typical male” (18–25 with prior-infection protection) will become COVID infected over the evaluation period if unvaccinated; e.g., 0.53 if $I_{r}=1$
According to Table 7, our modeling suggests that for a typical unvaccinated 18–25-year-old male with prior-infection, the choice to vaccinate created a hospitalization risk over twice that associated with the choice not to vaccinate even if $I_{r}=1$ (evaluation-period infection is certain for an infection-naïve male 18–25). Because a “typical male” has an “average comorbidities level,” the information in Table 7 is difficult to apply. However, if our typical unvaccinated male of Table 7 also has no comorbidities, then his hospitalization-risk ratios should exceed those of Table 7, [^12] See, e.g., the BMI 23–24 column of Table 9 because risk-ratio denominators should decrease given no comorbidities and there is no evidence that being free of comorbidities reduces VAM/P risk [67]. [^13] See also [68, p. 383].
@@
For an 18–25-year-old male without prior COVID infection, hospitalization-risk ratios become favorable for vaccination:
@@
Table 8. Hospitalization-risk ratios for a typical male 18—25 without prior-infection protection: risk of hospitalization after choosing mRNA-1273 vaccination (VAM/P or severe COVID-breakthrough infection) divided by risk of COVID hospitalization after choosing not to vaccinate.
@@
11 1400 in 1 million 323.4 in 1 million $0.42\approx\frac{268+323.4}{1400}$
@@ 4.2. Subpopulation Analyses Based on Prior-Infection Status and BMI
At this point, we discontinue confining to footnotes discussions and citations of data unavailable at the time of the FDA’s benefit-risk analysis of mRNA-1273.
The FDA acknowledges, “The health condition of individuals may remarkably impact the B-R profile and its evaluation” [11, Section 4]. Comorbidities such as obesity may triple COVID-hospitalization risk [69]. Our model can produce a benefit-risk analysis based on comorbidity status by appropriately adjusting its infection-hospitalization-rate input $H_{r}$. We illustrate the process in Appendix 3 of Supplement 1, focusing on obesity because evidence suggests that Body Mass Index (BMI) is the single most potent modifier of COVID-19 risk for those under 60 years old [70, 71]. We use an assessment of COVID-hospitalization risk based on BMI that finds risk is minimized when BMI $\approx 23$ and that for $23\leq\text{BMI}\leq 44$ risk is multiplied by a factor of approximately 1.09 (1.08, 1.10) for each unit increase in BMI [72]. The results of our modeling appear in Table 9 below.
Table 9. Hospitalization-risk ratios $\left(\frac{\text{risk if vaccinated (VAM/P \& breakthrough)}}{\text{risk if unvaccinated}}\right)$ for males 18–25 based on BMI . Ratios displayed as pairs: “with prior-infection protection, without prior-infection projection.”
@@
Consistent with Tables 7 and 8, Table 9 indicates that for those having BMI in the intermediate range 25–34, prior-infection status determines whether vaccination is beneficial. However, for the those with high BMI (35–44) in a high-infection-risk scenario Table 9 indicates that vaccination becomes beneficial even for those with prior-infection protection.
4.3. Validating the Model